OpenStudy (anonymous):

MEDALS!!!!! On a particular day, the wind added 3 km per hour to Alfonso’s rate when he was cycling with the wind and subtracted 3 km per hour from his rate on his return trip. Alfonso found that in the same amount of time he could cycle 36.5 km with the wind, he could go only 24 km against the wind. What is his normal bicycling speed with no wind? a. 15 km/h c. 35 km/h b. 40 km/h d. 45 km/h

OpenStudy (matt101):

Consider the distance (D), speed (S), and time (T) of each trip. Also call x the normal bicycling speed with no wind. For the trip with the wind, the distance traveled is 36.5 km, the speed is x + 3 km/hr, and the time is (x + 3)/36.5 hr (since T = D/S). For the trip against the wind, the distance traveled is 24 km, the speed is x - 3 km/hr, and the time is (x - 3)/24 hr. The question tells us the time for both trips is the same. In other words, (x + 3)/36.5 = (x - 3)/24. Solving for x you should get x = 14.52 km/hr which makes A) the correct answer.

OpenStudy (anonymous):

why is time for both trips equal

OpenStudy (anonymous):

@matt101 why did u consider the time equal for both trips?

OpenStudy (matt101):

It says it in the question: "Alfonso found that IN THE SAME AMOUNT OF TIME he could cycle 36.5 km with the wind, he could go only 24 km against the wind."

OpenStudy (anonymous):

oh yeah thnx